OpenStudy (anonymous):

On a survey at a particular university, 64% of the students indicated that they have misplaced their Student Identification Card at least once in the past year. If 10 students are in a campus computer lab, what is the probability that 2 or 3 of them have misplaced their Student Identification Card in the past year? Round to four decimal places, if necessary. A) 0.6442 B) 0.0299 C) 0.1583 D) 0.0236

OpenStudy (kropot72):

\[P(2\ students)=10C2\times 0.64^{2}\times 0.36^{8}\] \[P(3\ students)=10C3\times 0.64^{3}\times 0.36^{7}\] When you have calculated the above two values of probability, add the two values to find the required probability.